### v1

# BOUQuET [Uncaptioned image]: dataset, Benchmark and Open initiative for Universal Quality Evaluation in Translation

Abstract
This paper presents BOUQuET, a multicentric and multi-register/domain dataset and benchmark, and its broader collaborative extension initiative. This dataset is handcrafted in non-English languages first, each of these source languages being represented among the 23 languages commonly used by half of the world’s population and therefore having the potential to serve as pivot languages that will enable more accurate translations. The dataset is specially designed to avoid contamination and be multicentric, so as to enforce representation of multilingual language features. In addition, the dataset goes beyond the sentence level, as it is organized in paragraphs of various lengths. Compared with related machine translation (MT) datasets, we show that BOUQuET has a broader representation of domains while simplifying the translation task for non-experts. Therefore, BOUQuET is specially suitable for the open initiative and call for translation participation that we are launching to extend it to a multi-way parallel corpus to any written language.
1 Introduction
Although multilingual large-language model evaluation benchmarks are only starting (Dac Lai et al., 2023), there is a rich research history in multilingual evaluation datasets for natural language processing; e.g., Sun and Duh (2020); Malmasi et al. (2022); Yu et al. (2022), with Machine Translation (MT) being the task with the highest investment in multilinguality (Kocmi et al., 2024). This is evident from the nearly 15-year history of the Conference on Machine Translation (formerly a workshop, WMT), which has established an international evaluation campaign (Kocmi et al., 2024). The campaign has compiled a comprehensive collection of parallel data evaluations covering a broad range of language pairs, domains, and tasks. Although some datasets offer multi-way parallelism across languages (e.g., the TED parallel corpus), [^1] https://hal.science/hal-04702210v1 the largest multi-way parallel corpus was only recently introduced with FLORES-101 (Goyal et al., 2022), later expanded to FLORES-200 (NLLB-Team et al., 2022) and to 2M-FLORES (Costa-jussà et al., 2024).
However, even if great progress has been made, advances in multilingual applications for diverse languages are hindered by limited evaluations. Existing datasets and benchmarks fall short due to having an English-centric focus, a narrow register, compromised quality from automated construction and mining, limited language coverage, or a static nature, in addition to being prone to contamination.
Similarly, in parallel with the previous progress, there have been several initiatives that called for data annotation in a collaborative and open way, such as the translation data collection initiative reported in Singh et al. (2024).
This precisely sets the stage for the introduction of a multilingual evaluation dataset and a benchmark, BOUQuET, which additionally combines community efforts. The purpose of this paper is two-fold. First, the paper details how we develop the Source-BOUQuET dataset, which is the necessary stepping stone towards an open initiative. Second, it presents how we design the open initiative itself, which aims to build the Full-BOUQuET dataset; i.e., Source-BOUQuET translated into any written language.
1.1 Definitions
Before describing the Source-BOUQuET dataset’s characteristics and building methodology, we define our use of some frequently encountered terms that may cover a variety of meanings.
Domain.
By the term domain, we mean different spaces in which language is produced in speech, sign, or writing (e.g., books, social media, news, Wikipedia, organization websites, official documents, direct messaging, texting). In this paper, we focus solely on the written modality.
Register.
We understand the term register as a functional variety of language that includes socio-semiotic properties, as expressed in Halliday and Matthiessen (2004), or more simply as a “contextual style”, as presented in Labov , pp.79–99). In that regard, a register is a specific variety of language used to best fit a specific communicative purpose in a specific situation.
2 Dataset: Source-BOUQuET
In this section, we describe the creation criteria that have follow to design Source-BOUQuET, as well as the languages it includes.
2.1 Main characteristics
As described in greater detail next, the Source-BOUQuET dataset is mainly characterized by its non-English-centric focus, its diverse registers and domain (which is complementary to FLORES-200), its manual and original composition, and its built-in dynamic extensibility.
Non-English-centric focus.
Source-BOUQuET is handcrafted by proficient speakers of French, German, Hindi, Indonesian, Mandarin Chinese, Russian, and Spanish. Each of those languages provides the same number of sentences to the final dataset. The languages for Source-BOUQuET (see Table 2 in Section 2.3) are all part of the top 20 languages in the world in terms user population, as listed in Eberhard et al. (2024). In addition, they are also used by a large number of both native and non-native speakers, which makes them good candidates for what we refer to as pivot languages; i.e., higher-resource languages that can facilitate—as source languages—the translation of datasets into lower-resource languages. English is often used as such a pivot language, since numerous people have a high degree of proficiency in English as a second language. English is not the only language in this situation, however, and is not always the best pivot language option. For example, it is much easier to find Guarani-Spanish bilingual speakers than it is to find Guarani-English bilingual speakers. What is more, cultural proximity may also make translation slightly easier.
Diverse registers and domains.
Registers derive from communicative purposes and, as such, are related to domains. However, the relationship between registers and domains is not one to one. For example, if we take a domain such as TV news, we can identify at least 3 registers:
  • The register used by the news anchor, which is represented by fully scripted language that is read from a teleprompter with a very specific and unnatural form of diction (hypercorrect enunciation, unnatural intonation, homogeneous pace).
  • The register produced by communication specialists (i.e., people who have been trained to be spokespersons or surrogates). The points they make have been scriptedand rehearsed to the point of being known by heart. Communication specialists also learn rehearsed patterns of language to make sure they do not get surprised by a question and can redirect the conversation towards the points they need to make instead of answering unexpected questions. All the language they use has been written and rehearsed. It sounds spontaneous but it is not structured like informal language.
  • The register represented in person-in-the-street segments, which is more informal and spontaneous (possibly colloquial).
The example above is taken from a domain where both speech and writing are used but the situation is not significantly different in the written modality only. Language users all commonly shift between registers, which is typically referred to as style-shifting. Style-shifting (i.e., register-shifting) occurs within domains; so the domain itself is not a fool-proof way of getting a specific register. Although the norms of the domain can impose the degree of formality and of lexical specialization, it is often the register (which derives from the communicative purpose), not the domain, that determines many aspects of linguistic structure (lexical density, pronoun use, syntax, etc.).
The registers we cover is Source-BOUQuET are given in Figure 1.
Manual construction and original composition (not crawled) with accurate revisions
To develop Source-BOUQuET, we set a variety of linguistic criteria that need to be covered. Guidelines are then shared with linguists who manually craft sentences covering examples of these linguistic criteria and compose paragraphs of various lengths.
The main strategies for open collaboration are to design contribution guidelines and build the annotation tool that enables to freely collect translations in any language. For this, BOUQuET is initially composed on 8 major languages plus the English translation and it is shared in a repository that allows language communities to easily add a new language even if they do not speak English, e.g. a native speaker of Quechua can translate from Spanish. This repository contains detailed guidelines on how to do it. BOUQuET’s innovative approach ensures universal language accessibility. This open collaborative initiative will enrich BOUQuET with the following characteristics.
Language coverage extensibility
Using both private and community-driven initiatives, we could potentially support any written language, as long as there is individual interest in contributing to multilingual advancements.
Dynamic in nature
Since BOUQuET includes the community, it can continuously evolve by constantly engaging it.
2.2 Creation criteria
For the design of the creation guidelines, which are reported in detail in Appendix A, we prepared a list of linguistic coverage requirements as well as some statistical information.
Linguistic coverage requirements.
In order for BOUQuET to be representative of various linguistic phenomena, linguistic coverage requirements are defined (as listed in Table 1), which are to be included in sentences that form paragraphs. This continuity will include variation in the number of sentences per paragraph and in the lengths of sentences themselves.
Phenomena
Paragraph-like continuity
Variation in sentence lengths
Dominant (unmarked) and non-dominant (marked) word orders
Different emphasis or topicalization
Different sentence structures (affirmation, interrogation, negation, subordination, coordination)
Different verb moods, tenses, and aspects
Different morphosyntactic options
Different grammatical persons (1st, 2nd, 3rd, singular, plural)
Different grammatical genders
Different grammatical number agreement
Different grammatical case or forms of inflection
Most frequent words used in various registers
Presence of named entities, numbers, slang, and emojis
Table 1: Source-BOUQuET Linguistic Requirements
Variety of domains.
Source-BOUQuET is intended to cover 8 domains: narration (as in fiction writing), dialog, social media posts, social media comments, how-to, miscellaneous website content (excluding social media or news), opinion pieces, miscellaneous other (such as written speeches or signage). The choice of these domains optimizes for variety and popular usefulness.
Variety of registers.
Source-BOUQuET is built with register variety in mind, differently from FLORES-200, which covers a few different domains but remains largely within similar registers. We characterize the registers through 3 main aspects (connectedness, preparedness, and social differential), and a variety of options within these aspects (see Figure 1 for detail). Connectedness attempts to describe the type of interaction typically available in a given domain. Preparedness aims to gauge how much time is typically used to produce or edit language content. Social differential tries to describe various social situations in which language content is typically produced in each of the domains. Options are not mutually exclusive at the level of the domain but become differentiated at the level of a single sentence.
By including new registers and domains, the new dataset is likely to be more generalizable to different contexts and applications.
Figure 1: Covered register characteristics.
Statistical guidance.
In order to appropriately cover linguistic requirements and adequately represent the registers and domains, we performed a statistical analysis to understand each of the linguistic characteristics. In particular, our analysis covers most domains that we are including in Source-BOUQuET by using diverse public datasets: narration (Books3, gutenberg (Gerlach and Font-Clos, 2018)); Social media posts (Reddit (Baumgartner et al., 2020)); Social media comments (Wikipedia comments [^2] https://www.kaggle.com/competitions/jigsaw-multilingual-toxic-comment-classification); Conversations / Dialogues (dialogsum (Chen et al., 2021), Open Orca (Lian et al., 2023)); Tutorials/how-to articles (how-to Wikipedia-lingua  [^3] https://huggingface.co/datasets/GEM/wiki_lingua); Website content (c4 (Raffel et al., 2020)); News / Reflection pieces (cnndailymail (Nallapati et al., 2016), xsum (Narayan et al., 2018)) and Miscellaneous (Wikipedia). Note that we collect information from public data that do not always accurately match our categories but constitute a proxy. For each of these domains, we have analyzed dimensionality: characters per token; tokens per sentence and sentences per paragraph.
Figure 2: (Left) Tokens per sentence and (middle) sentences per paragraph (Right) CEFR per domain.
Regarding tokens per sentence (Figure 2 left), we can see correlations between different domains, and clear differences in length, especially in dialogs which tend to be much shorter.
Regarding sentences per paragraph (Figure 2 middle), we can find correlation between different datasets representing the same domain, where fiction writing paragraphs tend to be much longer (averaging 5 but reaching up to 20 ), dialogs and news articles are much shorter (barely reaching 3-4 sentences in a paragraph), and the rest of the categories are somewhere in between (normally staying between 1-5 but reaching up to 10 in some cases).
Furthermore, we have checked complexity: distribution of CEFR level [^4] https://rm.coe.int/common-european-framework-of-reference-for-languages-learning-teaching/16809ea0d4 as a proxy. Including here the % of C2 scores at sentence level for each dataset. This was labeled by a sonar-based model (Duquenne et al., 2023) trained on CEFR-SP data CEFR Text Classification which greatly outperformed Llama-3.1 (Touvron et al., 2023). See Figure 2 right.
Wikipedia seems to be the only dataset with a more considerable share of C2 sentences, with some others like dialogues having no samples scored as such.
Annotations.
Each entry of Source-BOUQuET includes the source text (in one of the 8 pivot languages from Table 2) and its translation into English, domain information, and contextual information for better translation accuracy.
2.3 Languages
BOUQuET aims to be as multicentric as possible, in contrast with most existing datasets like FLORES-200 that are English-centric. The motivation is mainly to be representative of linguistic phenomena. To this effect, it is created in 7 non-English languages. Each language contributes with a similar number of sentences along with their English equivalents given by the sentence creators themselves. See languages in Table 2.
ISO 639-3 ISO 15924 Language Family Subgroup1
cmn Hans Mandarin Chinese Sino-Tibetan Sinitic
deu Latn German Indo-European Germanic
fra Latn French Indo-European Italic
hin Deva Hindi Indo-European Indo-Aryan
ind Latn Indonesian Austronesian Malayo-Polynesian
rus Cyrl Russian Indo-European Balto-Slavic
spa Latn Spanish Indo-European Italic
Table 2: Source-BOUQuET Languages
3 Benchmark
Statistics
In total, BOUQuET currently contains 1,750 sentences. These sentences are split by doing a stratified selection among source languages and domains into development and test sets with an equivalent number of sentences. A portion of the development set (522) is intended to initially be kept hidden. Results in the following sections are presented with the test split of 889 sentences.
Table 3 provides a comparison of several relevant statistics from BOUQuET and the closest related datasets including FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (NLLB-Team et al., 2022).
Dataset Sent. Parag. Reg. Dom. Lang. Dyn.
FLORES-200 3,003 () $\times$× Wikipedia, News, Travel guides 200 $\times$×
NTREX-128 1,997 $\times$× News 128 $\times$×
NLLB-MD 9,000 $\times$× $\times$× Chat, News, Health 6 $\times$×
BOUQuET 1,750 Fiction, Conversation, Social media posts/comments 8+
Tutorials, Website, Reflection pieces, Miscellaneous
Table 3: Main statistics from MT evaluation datasets including BOUQuET: number of sentences, paragraph-information, registers, domains, languages, dynamism.
Domain representation
The performance of the model in a new or unseen dataset depends on the similarity between the dataset that was used to fit the model and the new dataset. In order to more broadly evaluate an MT model, we are interested in covering a different set of domains. We compare the domain coverage of BOUQuET with that of FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (Stolte et al., 2024) . To do this comparison, we represent data (multi-domains public datasets from Figure 2 in grey, each alternative dataset FLORES-200, NTREX-128 and NLLB-MD and BOUQuET) with SONAR (Duquenne et al., 2023). From SONAR vectors, we do a PCA-dimensionality reduction, see Figure 3. All domains from Figure 2 are represented in grey. Then, Figure 3 from left to right compares BOUQuET against FLORES-200, NTREX-128 and NLLB-MD, respectively. We qualitatively observe that BOUQuET covers a wider range of domains.
Additionally, to quantify this coverage, we measure the overlap between each dataset with each of the domains using two metrics, the Wasserstein distance (implemented with POT library [^5] https://pythonot.github.io/) and the Silhouette coefficient (implemented with sklearn [^6] https://scikit-learn.org/stable/). The Wasserstein Distance (WD), also known as the Earth Mover’s Distance (EMD), is a metric that measures the "effort" required to transform one probability distribution into another. Lower results indicate a higher similarity between clusters. The lowest consistent results for all domains are obtained with the data set BOUQuET. Alternatively, the Silhouette Score ($S$)  [^7] https://en.wikipedia.org/wiki/Silhouette_(clustering) measures how well each data point fits within its assigned cluster compared to other clusters, it ranges between -1 and 1 and the higher the coefficient, the more coherent the clusters are. Again, BOUQuET shows a lower score, consistent across domains, meaning closer representation across domains.
Figure 3: Domain representation and overlap with datasets (FLORES-200, NTREX-128, NLLB-MD)
Narration Website Opinion Dialog Social media Social media How-to Misc.
comments posts
WD SC WD SC WD SC WD SC WD SC WD SC WD SC WD SC
FLORES-200 0.32 0.05 0.32 0.03 0.32 0.02 0.31 0.11 0.32 0.09 0.31 0.08 0.31 0.06 0.33 0.03
NTREX-128 0.32 0.04 0.31 0.03 0.32 0.01 0.31 0.09 0.31 0.07 0.31 0.6 0.31 0.06 0.33 0.03
NLLB-MD 0.31 0.04 0.31 0.02 0.31 0.00 0.30 0.08 0.31 0.06 0.30 0.05 0.30 0.05 0.32 0.02
Bouquet 0.29 -0.02 0.29 -0.02 0.30 -0.03 0.28 0.02 0.29 0.01 0.28 0.00 0.28 0.00 0.30 -0.01
Table 4: Wasserstein Distance (WD) and Silhouette Score ($S$) for each domain and dataset. Lower WD and $S$ indicate better representation of the domain. Best results in bold.
MT Benchmarking
To help the reader understand why the dataset is useful, we present preliminary results to demonstrate its use for its intended purpose: MT benchmarking. We evaluated on translation systems (GPT-4o (Wu et al., 2024), Llama-3.1 (Llama-3.3-8B)(Touvron et al., 2023), and NLLB (NLLB-3.3B) (NLLB-Team et al., 2022)) using two metrics: Comet (CometKiwi-da-xl, range 0-1 and $\uparrow$ better) (Chimoto and Bassett, 2022) and MetricX (MetricX-24, range 0-25 and $\downarrow$ better) (Juraska et al., 2024), following WMT 2024 (Kocmi et al., 2024). See results in Table 5. We observe that BOUQuET scores consistently higher than NLLB-MD on average and for all models, suggesting that BOUQuET is easier to translate. This is an advantage for the open initiative, since the complexity of current MT test sets makes it harder to ask the community to participate in translations as it requires a high-level of expertise. Rankings across models nor languages is not preserved, which hints that BOUQuET may be posing different challenges to the models. We need to further investigate which linguistic challenges BOUQuET is adding. On this ranking, if we look at average BOUQuET, LlaMa is the worst system according to Comet but the best system according to Metricx. This does not happen in NLLB-MD, where the ranking is consistent across metrics. If we look per language, ranking across metrics is consistent for 4 out of 7 languages for NLLB-MD and for 1 out of 7 languages for BOUQuET. This raises a concern and may suggest that MT metrics themselves do not generalize well to the broader domains/registers in BOUQuET. However, this requires further investigation and we want to specifically explore different alternatives to use BOUQuET such as exploring the performance by domain and register. Note that we present only preliminary experiments, see Section 5 for more details on ongoing experiments.
M cmn deu fra hin ind rus spa AVG
COM($\uparrow$) MetX($\downarrow$) COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX
NLLB-MD
GPT4o 0.76 14.34 0.77 6.80 0.76 13.86 0.68 6.63 0.79 16.16 0.78 3.96 0.80 13.99 0.76 10.81
Llama 0.71 14.38 0.71 6.96 0.72 14.00 0.57 7.07 0.75 16.21 0.71 4.00 0.75 14.11 0.70 10.96
NLLB 0.61 14.80 0.70 7.01 0.70 13.98 0.64 6.72 0.73 16.20 0.72 4.43 0.76 14.06 0.69 11.03
BOUQuET
GPT4o 0.79 1.99 0.80 2.79 0.78 5.57 0.76 3.37 0.84 4.07 0.81 3.60 0.83 4.74 0.80 3.73
Llama 0.75 2.06 0.76 2.07 0.74 4.68 0.72 3.19 0.83 3.54 0.75 3.26 0.80 4.07 0.76 3.27
NLLB 0.67 2.19 0.79 2.44 0.77 5.30 0.74 3.49 0.82 3.94 0.78 3.40 0.67 4.42 0.77 3.60
Table 5: Results on Quality Estimation from English to 7 Source-BOUQuET languages on BOUQuET evaluation dataset .
4 Open initiative
Source-BOUQuET is intended to be translated into any written language. This ambition can only be achieved with the support of the community. For this, we have organized an open collaborative effort which involves language communities that are interested in contributing to this effort.
The purpose of this open initiative is to collect translations from Source-BOUQuET. To collect these annotations, we have chosen the ARGILLA tool [^8] https://argilla.io/ which is a collaboration platform for building datasets, designed for use by AI engineers and domain experts.
Together with setting the Source-BOUQuET in this tool, we have designed annotation guidelines which very much resemble those from FLORES-200 (NLLB-Team et al., 2022).
This open initiative is available in HuggingFace https://huggingface.co/spaces/facebook/bouquet and it is starting as part of the Meta Language Technology Partner Program [^9] https://docs.google.com/forms/d/e/1FAIpQLSdzcRdtkQCuTrXw727DgJgWbOPKDj5v0bArgGfQUTT6sEopFw/viewform publicly launched in AI Summit [^10] https://www.elysee.fr/en/sommet-pour-l-action-sur-l-ia.
5 Conclusions and Next Steps
In this paper, we have presented the Source-BOUQuET dataset and the attached open initiative. This dataset is available in 8 languages. We have shown that consistent BOUQuET gains in domain diversity by in two different metrics while keeping complexity lower than its competitors. The latter is specially relevant to simplify the translation for non-experts that may join the open initiative. This paper and Source-BOUQuET is only a fraction of the ambitious project that we are starting by launching this open call for community efforts. Please join us in making Universal Quality Evaluation in Translation available in any language!
Limitations
BOUQuET is work in progress; the BOUQuET dataset is still limited in number of languages and translations. The benchmarking is also limited in several axes (dataset comparison, models and metrics evaluated). However, we prioritized starting the BOUQuET open initiative in coordination with the Meta Language Technology Partner Program launch so that we could start raising the collaboration interest. As OmnilingualMT team, we continue actively investing on it, see next steps that we are planning to release soon.
Coming Soon
BOUQuET is actively evolving and we are currently working on making the following progress:
  • Adding 250 sentences in Source-BOUQuET in Egyptian Arabic.
  • Complete Source-BOUQuET in all 7 languages from Table 2 plus Egyptian Arabic.
  • The creation guidelines plus translation guidelines in all 8 pivot languages.
  • Designing quality control for each of the contributions and adding new languages to the incremental releases of BOUQuET.
  • Extending the Benchmarking by further showing the capabilities of BOUQuET, e.g. increasing the evaluation of linguistic signals over its alternatives (FLORES-200, NTREX-128, and NLLB-MD). For this, we propose to explore the performance in different domains and see if we get difference in the performance ranking of translation systems.
Acknowledgments
We would like to extend our sincere gratitude to Luis Cadavid, Radhi Datla, Suci Fitriany, Christine Haugen, Stephanie Mauro, Carissa Ikka Pardamean, Gauri Shanware, and Jyoti Singh for their invaluable contributions to the work presented in this paper.
Appendix A Specific Guidance for Paragraph and Sentence Creation
A.1 Overview
The Bouquet-source dataset needs to include 250 sentences in each of its source languages. This means that each linguist needs to create (i.e., create from scratch, not copy; see Section 2.4 above) 250 original sentences. These sentences must be:
  • Organized in logically structured paragraphs (see the Paragraphs section below)
  • Representative of the linguistic structures and features most frequently used in specific domains (see the Domains section below)
  • Representative of the most common register of language used in similar situations (see the Registers section below)
  • Accompanied by a gold-standard (i.e., best in class) human translation into English.
A.2 Paragraphs
The linguist will receive a template in the form of a spreadsheet, in which paragraph structures are designed and laid out. The template specifies the exact number of paragraphs and the exact number of sentences for each of the paragraphs. Each paragraph is given a unique paragraph ID (e.g., P01, P02, P15). Each sentence within each paragraph is also given a serial, non-unique ID (e.g., S1, S2, S3). The existing structure must not be modified in any way.
A.3 Domains
The template is divided into 8 domains:
  1. 1.
    How-to, written tutorials or instructions
  2. 2.
    Conversations (dialogues)
  3. 3.
    Narration (creative writing that doesn’t include dialogues)
  4. 4.
    Social media posts
  5. 5.
    Social media comments (reactive)
  6. 6.
    Other web content
  7. 7.
    Reflective piece
  8. 8.
    Miscellaneous (address to a nation, disaster response, etc.)
You must produce the set number of sentences for each of the domains; the domain / paragraph / sentence structure of the template cannot be changed.
A.4 Language Register Information
When creating sentences, make sure that the register of language being used is representative of the most expected and appropriate register for the situation. We understand that several registers could be possible; please use discretion when selecting a register, while making sure that the chosen register is among the most expected and appropriate. To help you make a determination, we define 3 main functional areas of language register:
  • Connectedness: What type of connection do language users who initiate the text have with other language users?
  • Preparedness: How much time do language users who initiate the text had or took to prepare the text?
  • Social differential: What is the relative social status of the language users who initiate the text towards other language users?
A.5 Linguistic Features
One of the main reasons for dividing the dataset into sections that correspond to domains is to attempt to cover as many registers and aspects of language as possible. For example, we know that:
  • Some pro-drop languages may drop the subject pronouns more often in some situations than in others.
  • Some case-marking languages may use some cases in specific situations but avoid them in others.
  • In English, lexical density increases when the level of formality increases.
  • Some languages use a specific past verb tense in storytelling, which stands out from other past verb tenses used in casual conversations or other situations.
  • Some languages use specific verb moods in some situations but avoid them in others.
A.6 Violating Content
While creating sentences, please be sure to avoid inserting violating content. Violating content is language that can fall under one (or more) of the below categories:
  • Toxicity
  • Illegal activities
  • Stereotypes and biases
A.7 Step-by-Step Description of Tasks
Please use the provided template to follow this step-by-step description.
Column A: Lang-ID This column should have the same 3-lowercase-letter code representing the source language of the sentences
being created followed by an underscore character ( _) and a 4-letter code representing the script.
Column B: Domain This is 1 of the 8 domains represented in the dataset (see Section 3.1).
Column C: Subdomain Please insert your description of the subdomain or topic.
Column D: P-ID This is the unique code identifying a paragraph (e.g., P01, P02, …, P58).
Column E: S-ID This is the non-unique code identifying the sequential place of the sentence within a paragraph.
Column F: Sentence In this cell, please type a sentence you created.
Column G: Translation into English After entering a sentence in your language in Column F, please provide a gold-standard human translation
in this cell.
Column H: S-Nchars This represents a count of the number of characters in the sentence.
Column I: S Comment_src_lang To help other linguists expand this dataset by translating your sentences into their own languages,
please add any comments that bring more context about the sentence.
Column J: S Comment_English Please provide an English translation of the comment your inserted in Column I.
Column K: Linguistic features Please list the register- or domain-specific linguistic features you tried to showcase in the sentence.
Column L: Connectedness Please use any of the options best describing the register area of Correctedness.
Column M: Preparedness Please use 1 of the options best describing the register area of Preparedness.
Column N: Social differential Please use any of the options best describing the register area of Social differential.
Column O: Formality Please indicate the level of formality best characterizing the sentence.
Column P: Relationship Please insert the intended relationship between the language users involved in the situation.
Column Q: Idea origin Please insert the name of the media type or platform that inspired the sentence.
Column R: P Comment_src_lang To help other linguists expand this dataset by translating your sentences into their own languages,
please add any comments that bring more context about the entire paragraph.
Column S: P Comment_English Please provide a translation into English for the comment you inserted in Column R.
Column T: P-Nchars This represents a count of the number of characters in the current paragraph.
Column U: Creator_Translator-ID Please insert your ID here, if it isn’t prepopulated.
A.8 Additional Guidance on Domain-Specific Content
Dialogues, especially those inserted in long creative writing (such as novels), often include the name of the speaker or a cue mark (e.g., — ), and sometimes quotation marks. When creating sentences for conversations, please feel free to invent names for speakers or to label speaker turns (e.g., A, B). Please place the names or speaker reference in markup tags, similarly to this: <Name:>or <A:>.
Emojis: There are emojis frequently in some social media and messaging domains. There should therefore be a representation, albeit very limited but the point isn’t to know how the system translates emojis because there are no real agreed ways to translate them across hundreds of languages.
Social media comments: Okay if the structure is flattened. Okay to have some tags but not absolutely necessary and no need to insert them everywhere.
Disfluencies in informal conversations?: If they are representative of conversations and they can be translated (i.e., there is some consensus on how to write them in the language — ah, oh, um).
Appendix B Contribution Statement
Pierre Andrews: Open initiative; Supervision/Leadership; core; tool deployment and design
Mikel Artetxe: Benchmarking; Supervision; support
Mariano Coria Meglioli: Dataset; core; statistical analysis
Marta R. Costa-jussa: Dataset, Benchmarking, Tooling Supervision/Leadership; core; corresponding author; research direction, paper editor
Joe Chuang: Dataset; core; source-bouquet creator
David Dale: Benchmarking; Supervision/Leadership; core; evaluation and experimental design
Cynthia Gao: Dataset; support
Jean Maillard: Open initiative; core, licensing
Alex Mourachko: Management; support
Christophe Ropers: Dataset and Open initiative; Supervision/Leadership; core; dataset design and creation, guidelines design and creation, paper editor
Safiyyah Saleem: Management; support
Eduardo Sánchez: Benchmarking; core; experiments deployment
Yiannis Tsiamas: Benchmarking; support
Arina Turkatenko: Dataset; core; dataset creation
Albert Ventayol-Boada: Dataset; core; dataset creation
Shireen Yates: Management; support